solve: (2.4)^x = (2.6)^x - 1.. plzz provide full soln.....

solve:
(2.4)^x = (2.6)^x - 1..

plzz provide full soln.....

Grade:12

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
Please find the answer to your question below
Given(2.4)^x = (2.6)^x - 1
So (24)x=(26)x-(10)x
By interchanging the terms we get
(26)x-(24)x=(10)x
By substitution method by putting x=1 doesnot satisfy the equation
By putting x=2 we get LHS=262-242=676-576=100
RHS=102=100
Hence x=2 is the solution